Table 1. The estimated rate of E̸_{T} + ≥ 2 jet events from the largest standard model sources.

Background Source Number of Events
Z → νν̄  23 ± 8
W → τν̄  41 ± 15
W → μν̄  18 ± 6
W → eν̄  9 ± 3
Heavy Quarks  4 ± 4

the largest being the production and leptonic decay of W bosons and the production of Z bosons following by the decay into neutrino pairs. These background rates have been estimated using the measured boson cross sections and the ISAJET Monte Carlo programme^{15} to model the kinematics of the W and Z boson daughters. The other significant potential source of large E̸_{T} events are from pair production of top or bottom quarks. These rates have been estimated assuming a top quark with a mass of 91 GeV/c^{2}, which gives the most conservative estimate for this background. The estimated background event rates are listed in Table 1. The total background estimate is 91 ± 17 events, which agrees well with the observed rate.

A limit on the pair production of q̃’s and g̃’s is derived by fitting the observed E̸_{T} distribution to contributions from the background sources and from SUSY production for various choices of Mq̃ and Mg̃. Both the fraction of SUSY and background events are allowed to vary in the fit. The likelihood distribution for the number of SUSY events in the data is then used to set an upper limit on the SUSY pair production cross section as a function of the squark and gluino mass. The region in the Mq̃-Mg̃ plane where this upper limit is below the predicted cross section is shown in Fig. 5. The mass limits obtained when either the q̃ or g̃ is not present (or very massive) are

Mg̃ > 141 GeV/c^{2} at 90% C.L.  (8)
Mq̃ > 126 GeV/c^{2} at 90% C.L.  (9)

It has been pointed out that for such large q̃ and g̃ masses, the simple decay scenario assumed in establishing these limits is unlikely.^{16} More complex “cascade” decays of squarks and gluinos are expected involving additional supersymmetric particles that change the expected signature by increasing the number of daughter particles and softening their P_{T} spectra. This would tend to reduce the efficiency of the event selection and accordingly lower the mass bounds shown in Fig. 5(a). The size of this effect depends on several parameters of the model making it difficult to summarize the mass limits; however, the mass limits are typically reduced by 20-30 GeV/c^{2} for most parameter choices. The excluded region of squark and gluino masses for a specific set